OpenStudy (tylerascough):

Need help will fan and medal Express 31/10 as a decimal

OpenStudy (mahoganie.carson):

Change 3 1/10 as an improper fraction. 3*10 = 30 30+1 = 31 31/10 3 1/10 as an improper fraction is 31/10. Convert 31/10 to a decimal by dividing the numerator by the denominator. 31/10 = 3.1 3 1/10 as an improper fraction is 31/10 as a decimal is 3.1. 3 1/10 as a decimal is 3.1.

OpenStudy (tylerascough):

Thank you so much mahoganie.carson

OpenStudy (kittiwitti1):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TylerAscough Need help will fan and medal Express 31/10 as a decimal \(\color{#0cbb34}{\text{End of Quote}}\) Please close the question if you are done. :)